adding and subtracting rational expressions

3/(2x) + 5/(3y)
1
Expert's answer
2012-12-03T09:04:14-0500
3/(2x) + 5/(3y)=(3*3y)/(2x*3y)+(5*2x)/(3y*2x)=(9y+10x)/(6xy)